OpenStudy (anonymous):

I have a math final today and i need to know more about solving a quadratic equation by completing the square. A sample problem is d^2+3d-10=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

1/2 coefficient of d; square it; add it to both sides

OpenStudy (anonymous):

the two roots of a quadratic equation of the the type \[ax^2 +bx +c=0\] are \[-b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (anonymous):

Hi Dimitri, Solution is as follows

OpenStudy (anonymous):

first u shift the constant ie -10 to RHS to get d^2 + 3d = 10

OpenStudy (anonymous):

\[-3\pm \sqrt{9--40}/2\] hence the solutions are (-3+7)/2 =2 and (-3-7)/2 =-5

OpenStudy (anonymous):

Now add square of half of coefficient of d i.e. (3/2)^2 to both sides to get d^2 + 3d + (3/2)^2 = 10 + (3/2)^2

OpenStudy (anonymous):

LHS becomes equal to (d + 3/2)^2, so we get (d + 3/2)^2 = 10 + 9/4 Hence (d + 3/2)^2 = 49/4

OpenStudy (anonymous):

Further taking root of both sides, we get d + 3/2 = root(49/4) so d + 3/2 = +/- 7/2 d = +/- 7/2 - 3/2

OpenStudy (anonymous):

So u get two values of d d = 7/2 - 3/2 = 4/2 = 2 and d = -7/2 - 3/2 = -10/2 = -5

OpenStudy (anonymous):

Is that clear for u Dimitri????